# TermSight: Making Service Contracts Approachable

### v1

@@ 7.2.2. Summary driven reading
19 out of 20 participants at times relied on the summary snippets as a primary entry point for engaging with the text. One notable example of this behavior was P3, who mainly read the Summary Snippets and clicked 11 of them to interrogate the original text (Figure  14a). P3 described how colors were driving his attention to read. P3’s decision to click the Summary Snippets and interrogate the original text relied on the color coding and evaluation of the information loss of the Summary Snippets.
@@
7.2.3. Original text driven reading
At other times participants centered their interactions around the original text. 9 participants went through at least one policy focusing mainly on the original text. These participants described how they wanted to first read the actual text before relying on AI-generated summaries due to the potential imperfections of AI. These participants explained, “I didn’t want my opinion to be affected by AI” (P5) and “sometimes summaries don’t have all of the information from my other experiences” (P18).
When reading the Privacy Policy, P8 focused on the original text as shown by the frequent scrolling of the original text in Figure  14b. In addition, P8 clicked the Highlight Bar 11 times to refer back to the Summary Snippets. P8 described how he knew AI was not perfect from his prior professional experiences in AI. As a result, P8 took the colors and summary snippets as a guide rather than a guarantee.
@@ 7.2.3. Original text driven reading
7.2.4. Calibration of trust
6 participants described that they went through a calibration process with AI-generated features such as the Summary and Color at the start of the reading session. For example, P6 described how his reading behavior was different during the calibration process and after. P6 described reading the original text on the right side to evaluate the colors and summaries at the beginning of the reading session (original text driven reading). Afterward, P6 relied on the summaries a lot more by scrolling and skimming the left panel and clicking on the summary snippets to dive to the original text as seen in Figure  14c (summary driven reading).
“So like early on in the reading, I kind of read the whole thing on the right first. Then, I looked at the summaries, and once I realized that it did do a good job summarizing it with colors. I trusted it enough to trust it for the rest of the time.” (P6)
7.2.5. Switching strategy based on time and document.
ToS includes multiple documents, and in the study, there were 14-15 policies with varying lengths. We noticed how participants’ reading behavior can vary based on document length (P5, P19). For example, P19 mainly read the summary for longer policies such as User Agreement and Privacy Policy (Summary driven reading). For shorter policies such as Buyer Policy, P19 only read the original text (Original text driven reading).
We also noticed how time can change participants’ reading strategy (P7, P8). For example, P7 spent 8.5 minutes and was very interrogative when reading the User Agreement policy. When she realized that she was running out of time, she skimmed through the summary snippets for 5 more policies without interrogating the original text.
7.3. RQ3: How did TermSight influence comprehension and recall?
Below we present the results of the comprehension quiz (§7.3.1) and recall task (§7.3.2).
7.3.1. Comprehension
For the comprehension quiz, out of 6 questions, participants on average scored 2.00 ($\sigma=1.03$) when using TermSight and 2.15 ($\sigma=1.23$) when using the baseline interface. Our Bayesian model analysis for the comprehension quiz (model details in Appendix  D.2) reveals no significant differences in the comprehension scores across interface conditions. As shown in Figure 15, the posterior distributions for both conditions were nearly identical and centered around zero. When the question ID and service type are fixed, there are significant overlaps between the 94% HPDI for each treatment–control pair as shown by the forest plots in Figure 20 in the Appendix, suggesting no significant differences.
Figure 15. The figure shows the posterior distributions of the coefficients for the treatment and control conditions. Both posterior distributions are nearly evenly split around zero, indicating no effect of the treatment on the comprehension quiz questions.
7.3.2. Recall
For the recall task, participants using TermSight recalled 39 correct facts and 3 incorrect facts in total. On the other hand, participants using the baseline interface totally recalled 26 correct facts and 7 incorrect facts. Our Bayesian model analysis for the recall task (model details in § D.3) reveals no significant differences in the recall scores across interface conditions. Figure 16(b) shows the posterior distribution of the effect size of the treatment on recall ($\mu=0.6$, HPDI = [-0.39, 1.6]). Despite the mean being 0.6, suggesting a medium effect size, the HPDI interval has an overlap with the ROPE, implying that there is no significant effect of the treatment on the user recall.
(a) Posterior distributions of the coefficients for the treatment and control conditions.
(b) Posterior distribution of the effect size of the treatment on user recall.
Figure 16. Figure 16(a) shows the posterior distributions of the coefficients for the treatment and control conditions. Specifically, the posterior distribution of the treatment condition is shifted to the right, indicating a positive influence of the treatment on Recall. Figure 16(b) shows the posterior distribution of the effect size of the treatment on the user recall centered around 0.6, suggesting a medium effect size. Yet, the HPDI has an overlap with the ROPE of [-0.1, +0.1], implying no significant effect of the treatment on the user recall.
@@ 8. Discussion
In this paper, we set out to design a new reading experience that scaffolds readers’ ability to find and understand information important to them within Terms of Service (ToS) contracts. Based on formative interviews and observations, we realized this vision in TermSight. Results from our evaluation of TermSight (N=20) suggest that TermSight made it easier to read, navigate, and contextualize information within ToS relevant to users. Especially interesting was how TermSight afforded multiple different reading strategies (e.g., summary-driven reading, §7.2.2), all of which helped participants overcome the key difficulties of contract reading we observed in our formative study (§3). However, similar to prior works on simplifying ToS or other complex documents  (Vu et al., 2007; Tabassum et al., 2018; Good et al., 2005; Korunovska et al., 2020; Robinson and Zhu, 2020; August et al., 2023), TermSight did not worsen or improve the comprehension or recall of the content. Below, we discuss the implications of our findings and the TermSight system.
8.1. Existing ToS are not designed to be read
Many existing ToS have static or interactive design elements, such as policy names, section headers, and overview summaries (Vu et al., 2007). However, participants in our formative study found these features vague, misleading, and ineffective for surfacing and interpreting relevant information (§3.3). The HCI literature often refers to these ineffective and deceptive designs as dark patterns in the user interface that manipulate behavior or obscure information to serve the interests of the service providers  (Mathur et al., 2021; Bösch et al., 2016; Kyi et al., 2023).
TermSight offers one possible design to overcome these dark patterns by offering contract-level guidance (e.g., Power Meter and document previews), document-level guidance (e.g., color-coded Summary Snippets), and phrase-level guidance (e.g., phrase identification). Participants in our usability study (§7) affirmed these design goals in both self-report and behavior metrics. Participants reported that features of TermSight helped them decide which policy to read (Power Meter, Document Preview) and what text to read within a policy (Summary Snippet). Moreover, participants highlighted how TermSight made the original text more approachable with Summary Snippets that broke down and simplified dense sections and Phrase Scope that clarified unfamiliar phrases and helped envision implications.
Future systems could take these designs further by adapting them to other types of contracts, such as leasing contracts, medical contracts, or employment contracts. For example, Summary Snippets may help readers surface and interpret relevant or predatory clauses in leases, while phrase scope may help patients envision the implications of clauses in the medical contract through patient-specific scenarios. Lastly, while we designed TermSight as a contract reading interface, TermSight can be readily extended as a browser extension that highlights relevant and important clauses based on user-defined personas to provide in-situ legal awareness and warnings when encountering contracts.
8.2. Envisioning consequences to legal text
Prior work has explored providing term definitions to aid comprehension of academic text (August et al., 2023; Head et al., 2021). Yet, participants in our study mentioned that definitions alone were not always sufficient for legal language. Participants noted that it can be difficult to contextualize abstract phrases in real-world scenarios or anticipate potential consequences—especially those they have not yet personally experienced. Our formative study further revealed how people’s intended usage of the service, personal values, and worries can vary. While prior work explored predicting unintended consequences for scientific endeavors and AI  (Wang et al., 2024; Pang et al., 2024), TermSight explored the potential value of presenting personalized scenarios to illustrate the possible implications of legal phrases before a user agrees to a ToS. Some participants even noted how the scenario was more helpful than the definition (§7.1.5). By grounding abstract terms in concrete, relatable examples, these scenarios may help users better reflect on the potential implications behind legal language. Future work could investigate additional ways to help consumers reflect on unintended consequences, such as by showing multiple scenarios or relevant legal cases.
8.3. Design for calibration and trust
LLM output can be imperfect due to hallucinations or confusing output  (Huang et al., 2024; Maynez et al., 2020; Luo et al., 2024). It is therefore important to help users form appropriate trust and reliance for fostering effective Human-AI collaboration  (Schemmer et al., 2023; Cao et al., 2024). In this vein, prior reading interfaces integrating AI have introduced cognitive forcing functions (Buçinca et al., 2021), such as showing summaries only on demand (August et al., 2023) or at the end of the reading session (Chen et al., 2023), to encourage readers to first read the original document. Yet, it is unclear if these designs are always optimal for intelligent reading interfaces, since they can increase user frustration and may limit specific reading behaviors  (Buçinca et al., 2021). For example, these forcing functions, if applied to TermSight, might introduce significant friction for participants taking a summary-driven reading approach (§7.2.2).
TermSight took the approach of tightly linking the AI-generated summaries to the original text to allow users to compare and contrast the two. In TermSight, users can navigate between and compare the summary and the original text in one click. When given these affordances, participants took advantage of them. Some participants explicitly evaluated the Summary Snippets to calibrate their trust. Interestingly, one participant framed the evaluation of AI-generated output as his ‘due diligence’. Even those who took a summary-driven reading approach frequently interrogated the referenced original text as opposed to solely reading the summaries. Future work can explore how to help users calibrate trust with the system and evaluate AI output. This could include providing explanations of how the reading interface works in general, providing examples of AI imperfections, and providing explanations for individual features  (Schemmer et al., 2023).
8.4. Ethical Considerations and Social Implications
Features of TermSight (e.g., Summary Snippets, Classification of Power and Relevance, Term Definitions, and Scenarios) rely on the capabilities of LLMs. However, LLMs have been shown to produce imperfect outputs and factually incorrect hallucinations  (Huang et al., 2024; Maynez et al., 2020; Luo et al., 2024). This introduces risks when interacting with LLM-generated text, as they are not part of the contract and are not legally binding. For example, incorrect summaries might cause readers to misinterpret the clauses in the contract that are binding. On one hand, LLMs have already shown strong capabilities in generating summaries (Zhang et al., 2024; Liu et al., 2024b), term definitions (Yin and Skiena, 2023; Jhirad et al., 2023), and scenarios (Jiang et al., 2024; Wang et al., 2024; Feng et al., 2024; Shaikh et al., 2024). Work is also actively investigating how to detect non-factual LLM output  (He et al., 2024) and how to guide LLMs to generate factually correct information  (Vu et al., 2024). As a result, LLM output would ideally become more factually correct over time. On the other hand, for future deployment of systems like TermSight, we still believe a validation mechanism for LLM output is necessary. For example, future implementation of TermSight’s features (e.g., term definitions and scenarios) can be accompanied by additional sources of information when available (e.g., news articles, court cases, or information from other ToS). Furthermore, TermSight could facilitate end-user auditing (Lam et al., 2022; Jahanbakhsh et al., 2022), allowing users to report inaccuracies. These user contributions could collectively help validate and refine model outputs.
@@ 8.5. Limitations and Future work
The generalizability of TermSight and our study findings is limited by the challenges associated with handling ToS documents and our sample of participants in the user study. TermSight currently relies on a set of precleaned Markdown or HTML files. Future works can investigate more advanced scraping and cleaning methods. In addition, our recruited participants were mostly college-educated and fluent in English. Marginalized communities, such as neurodivergent readers and individuals with limited English literacy, can face additional barriers when interpreting legal documents. For example, readers with ADHD might have a hard time staying focused when reading visually dense text such as ToS  (Stern and Shalev, 2013; Bental and Tirosh, 2007). Future works can investigate how to help marginalized populations make sense of legal contracts by evaluating TermSight with these populations of users or integrating additional supports such as translation.
@@ 9. Conclusion
ToS are service contracts outlining the rules and expectations for both the service provider and the consumer. Our formative study revealed ineffective and deceptive designs at all granularities of a ToS. To make ToS approachable for consumers, we designed and evaluated an intelligent reading interface: TermSight. Participants reported that TermSight reduced the difficulty of engaging with ToS and increased their willingness to do so. Participants highlighted that features of TermSight afforded them the ability to surface and interpret relevant information at all levels of ToS reading. They found the Summary Snippets with colors particularly helpful in breaking dense sections into snippets of simplified information, which helped them surface relevant information and power balance. The ‘duty to read’ originated when contracts were brief and based on mutual understanding and negotiation  (Kar and Radin, 2019; Benoliel and Becher, 2019). Modern ToS contracts, however, prioritize unilateral control over mutual agreement by design. TermSight presents a vision and serves as a technology probe into making legal contracts more accessible to the general public, providing consumers with the tools to make more informed legal decisions.
Appendix A Iterative Design
A total of 8 participants were recruited through Prolific to evaluate an early prototype of TermSight. This prototype closely resembled the current version of TermSight but lacked a document preview when hovering over the Power Meter. Additionally, the prototype used three saturation levels for relevance, resulting in a total of nine colors (three colors for power X three saturation levels for relevance). Participants could also toggle between two layouts for the Summary Snippets: condensed layout and in-context layout. The condensed layout matched the current design shown in Figure  1. In contrast, the in-context layout placed each Summary Snippet directly next to its corresponding Information Snippet. After a brief interface tutorial, participants were given 10 minutes to read a ToS using the interface followed by a semi-structured interview about their experience. Each session lasted for 40 minutes. Below we present the main findings and changes made.
General findings.
All participants described that reading with the prototype was easier and more engaging compared to typical ToS, which they tend to skip entirely or skim superficially. In contrast to their usual approach of skimming and trying to search for keywords that catch their eyes, all participants described how the Summary Snippets directly highlighted relevant information otherwise buried in dense sections. Additionally, participants expressed interest in reading the original text after using the Summary Snippets to surface relevant information and valued the ability to navigate between the summary and the original text. 6 participants found the scenarios provided by Phrase Scope to be particularly helpful as the definitions alone can be difficult to contextualize and envision implications.

### v2

@@ 7.3.2. Reading Strategies for Accessing Original Text
19 out of 20 participants at times relied on the Summary Snippets as a primary entry point for engaging with the text. One notable example of this behavior was P3, who mainly read the Summary Snippets and clicked 11 of them to interrogate the original text (Figure  12a). P3 described how colors were driving his attention to read. P3’s decision to click the Summary Snippets and interrogate the original text relied on the color coding and evaluation of the information loss of the Summary Snippets.
@@
Original text driven reading
On the other hand, nine participants went through at least one policy focusing mainly on the original text. These participants described how they wanted to first read the actual text before relying on AI-generated summaries due to worries about the potential imperfections of AI. When reading the privacy policy, P8 focused on the original text as shown by the frequent scrolling of the original text in Figure  12b. In addition, P8 clicked the Highlight Bar 11 times to refer back to the Summary Snippets. P8 described how he knew AI was not perfect from his prior professional experiences in AI. As a result, P8 took the colors and Summary Snippets as a guide rather than a guarantee.
@@
Calibration of trust
Six participants described that they went through a calibration process with AI-generated features such as the Summary Snippets and color coding. For example, P6 described how his reading behavior was different during the calibration process and after. At the beginning of the reading session, P6 focused on reading the original text to evaluate the colors and summaries (original text driven reading). Afterward, P6 relied on the summaries a lot more by scrolling and skimming the left panel and clicking on the Summary Snippets to dive to the original text, as seen in Figure  12c (summary driven reading).
“Early on in the reading, I read the whole thing on the right first. Then, I looked at the summaries, and once I realized that it did do a good job summarizing it with colors. I trusted it enough for the rest of the time.” (P6)
@@ 8. Discussion
In this paper, we explored how AI-powered systems might support ToS contract reading given the unique challenges that legal text poses (i.e., value-dependent, ambiguous, and legally binding). Our formative study revealed the barriers ToS presents, which informed the design of an augmented reading interface, TermSight. Results from our user study (N=20) suggest that TermSight’s features made ToS contracts easier to read and navigate, without an increase or decrease in comprehension. Below, we discuss the implications of our findings around the barriers ToS contracts present (§8.1), the design implications for future contract reading interfaces (§8.2), and the limitations of technology in mitigating the societal difficulties with service contracts (§8.3).
8.1. From Simplifying Legal Text to Navigating Dark Patterns
Past work on augmenting legal text has focused on simplification (e.g.,  (Manor and Li, 2019)). In the context of ToS contracts, our formative study revealed that the barriers people face extend well beyond complex text. Participants struggled to navigate multiple policies within a ToS, surface relevant information from dense legal language, and interpret ambiguous terminology. Surprisingly, existing navigational affordances—such as policy names, section headers, or overview summaries—often obfuscated rather than clarified the original text, a design reminiscent of dark patterns in the HCI literature  (Mathur et al., 2021; Bösch et al., 2016; Kyi et al., 2023). These barriers may not be unique to ToS, but reflect broader challenges for reading legal contracts taken to their extreme. For example, leases, insurance policies, and employment contracts similarly feature multiple addenda; long, dense text; and ambiguous language that obscures consumer rights  (Mueller, 1970; Furth-Matzkin, 2017). Future work could expand our findings by conducting large-scale analyses of dark patterns in contracts, an area largely underexplored by the HCI community. In addition, future contract reading interfaces should consider going beyond simplifying legal text to helping users navigate contracts while combating the dark patterns of design. TermSight serves as an initial exploration for inspiring future contract reading tools in this direction.
8.2. Design Implications and Opportunities
TermSight is a preliminary exploration of designing contract reading interfaces. Below, we discuss the design implications of our findings and opportunities for further exploration.
8.2.1. Providing guidance at multiple levels of granularity
While prior work focuses on simplifying legal text (e.g.,  (Manor and Li, 2019)), TermSight demonstrated the potential values of guiding readers at the contract-level (e.g., Power Meter and document previews), document-level (e.g., color-coded Summary Snippets), and phrase-level (e.g., phrase identification). Participants in our user study affirmed these designs in both self-report and interaction behaviors (§7). For example, participants reported how Power Meter transformed deciding which policy to read from a “guessing game” to a more deliberate decision-making process. On the other hand, Summary Snippets with color and Phrase Scope drew participants’ attention to snippets of information and phrases buried in dense text that they might otherwise miss. Consequently, finding which policy to read and what text to read within a policy was significantly easier for participants with TermSight. The specificity of TermSight’s guidance (i.e., relevance classification) relies on the user persona. For the user study, we designed two personas to be able to control for their quality (§6.1.2). These personas also aligned with participants’ profiles in the user study (Figure  13). Future work could explore allowing users to further customize their personas to receive personalized guidance and overcome the barriers of doing so (e.g., forming information bubbles  (Sharma et al., 2024)).
8.2.2. Prompting reflection on unintended consequences
Prior work has explored providing term definitions to aid reading academic papers (August et al., 2023; Head et al., 2021). However, legal contracts are unique because contracts are agreements for the unforeseen future rather than a description of the past  (Hart, 2017). Reflecting this characteristic of contract, participants in our formative study mentioned that definitions alone may not always be sufficient for legal language because it can be difficult to contextualize abstract phrases in unforeseen future scenarios. In the user study, participants described the value of Phrase Scope for presenting scenarios to prompt reflection on possible implications of signing a contract, similar to how prior work has encouraged scientists to reflect on the unintended consequences of their research (Wang et al., 2024; Pang et al., 2024). Some participants even noted how the scenario was more helpful than the definition (§7.1.5). Future work could further explore how to design these scenarios and investigate their influence on user perception of the contractual clauses, such as by manipulating the linguistic framing (Sharma et al., 2023) or showing information from legal cases and news articles.
8.2.3. Centering augmentations around the original text
Legal contracts differ fundamentally from documents like academic papers or news articles because the language itself carries legal power (Chen-Wishart, 2012). Consequently, LM-powered text transformations cannot substitute the original text. TermSight explored the approach of tightly linking generated text with the original text through visual placement (i.e, side-by-side views) and deep linking (i.e., linking to narrowly attributed text like a single clause), allowing readers to directly compare language and navigate to the original text in at most one click. When given these features, participants took advantage of them to engage with the original text (§7.3.2). Some participants dove into the original text by finding relevant Summary Snippets, while others verified their understanding of the original text with the Summary Snippets. Some participants even explicitly evaluated the Summary Snippets with the original text to calibrate trust in the system. An exciting avenue for future work would be to go beyond supporting original text reading to encouraging it. Features that nudge readers toward the original text where appropriate, such as revealing summaries only after reading  (Chen et al., 2023), are an interesting start, but little work has evaluated these features’ effects in situations where the original text might be discouraging to engage with.
8.3. Technological Limitations and Legal Implications
Compared to simplified reading tasks (e.g., reading a single clause (Robinson and Zhu, 2020; Lee et al., 2025) or policy overview (Tabassum et al., 2018; Korunovska et al., 2020)), our user study involved realistic ToS with over 10 policies. TermSight made ToS significantly easier to read and navigate. However, TermSight did not worsen or improve comprehension, reminiscent of prior findings that simplified language often did not improve comprehension of privacy policies  (Tabassum et al., 2018; Korunovska et al., 2020; Robinson and Zhu, 2020). There may be multiple explanations behind this finding. For example, Summary Snippets may have distracted attention away from the original text, the 12-word limit of Summary Snippets may have constrained the conceptual coverage of the original text, and participants’ calibration behaviors (i.e., confirming summary information with the original text) may have distracted them from the reading task. However, our findings may have also revealed the broader limitations of technological solutions in mitigating the societal difficulties with service contracts such as ToS. Although TermSight offered multi-level guidance on relevance (e.g., Power Meter, Summary Snippets), the sheer volume of potentially relevant information in ToS remains a dominant challenge. For example, in the social media ToS, 172/759 (23%) discrete Information Snippets were classified by TermSight as being relevant to the given user persona (§6.1.2), including 6 snippets needed to answer the comprehension questions. Within this subset, being able to notice and recall snippets of information to correctly answer the comprehension questions may be difficult. While future deployment of TermSight could allow users to customize the personas for more targeted guidance, this may risk introducing information bubbles  (Adamic and Glance, 2005; Conover et al., 2021; Garrett, 2009; Sharma et al., 2024). Given that the formation of a binding contract under U.S. law is premised on the doctrine of the ‘duty to read’ [^6] Contracting parties are legally responsible for reading and understanding the contract before signing (Benoliel and Becher, 2019)., our findings raise fundamental questions about the practicality of this duty for ToS. Instead of relying solely on technological fixes, policy and legal innovations are needed. For example, policy makers and legal scholars could explore alternative paradigms of contracting that operationalize ToS into shorter, context-specific contracts to reduce the upfront burden to consent.
@@ 8.4. Limitations and Future Work
The nature of this work is exploratory. Similar to prior HCI studies (e.g.,  (Chen et al., 2023; Cao et al., 2023; Kambhamettu et al., 2024)), we designed and studied TermSight to explore the challenges and opportunities of designing contract reading interfaces, rather than proposing TermSight as the definitive solution. Future work can pursue finer-grained exploration through controlled ablation studies. Examples may include investigating the influence of Power Meter on ToS reading frequency in the wild, assessing how hypothetical scenarios affect user perception of clauses (e.g., trust  (Shneiderman, 2000)), or investigating how AI affordances (e.g., Summary Snippets) support or distract from original text reading (e.g, (Nielsen et al., 2023)) through eye-tracking.
Because participants were paid to read ToS in a timed reading task, their motivation and behavior may not reflect real-world usage. Moreover, a within-subject design may lead to ordering effects or fatigue. In the study, we counterbalanced the conditions and didn’t observe a significant effect of ordering on user experience, comprehension, and recall (Appendix Figure  16). Yet, it may still have influences on participants’ interaction. Future work can expand our study by conducting between-subject or field experiments.
In addition, our participants were recruited from Prolific, mostly college-educated, and fluent in English, which may limit the generalization of our findings to the broader population. Marginalized communities, including neurodivergent readers and individuals with limited English literacy, can face additional barriers when interpreting legal documents. For example, readers with ADHD might have a hard time staying focused when reading visually dense text such as ToS  (Stern and Shalev, 2013; Bental and Tirosh, 2007). Future work can investigate how to help alternative reader populations make sense of legal contracts by evaluating TermSight with these populations of users or integrating additional supports (e.g., translation).
Future systems could adapt our design of TermSight to other contracts. For example, Summary Snippets may help readers surface and interpret relevant or predatory clauses in leases, while Phrase Scope may help patients envision the unintended consequences of signing a medical contract through patient-specific scenarios. Finally, features of TermSight rely on the capabilities of LMs, which may produce imperfect outputs  (Huang et al., 2024; Maynez et al., 2020; Luo et al., 2024). While work is actively investigating how to guide LMs to generate factually correct information  (Vu et al., 2024), we believe a validation mechanism for LLM output is necessary for future deployment of systems like TermSight, such as by facilitating end-user (Lam et al., 2022; Jahanbakhsh et al., 2022) or expert auditing of AI output (Jiang et al., 2024).
@@ 9. Conclusion
We regularly sign legal contracts for where we live, who we work with, and how we interact with digital services. Yet, these contracts have drifted from their premise of facilitating mutual understanding to complex documents that discourage reading. In this work, we explored the opportunities and limitations of intelligent reading support for legal contracts, using ToS contracts as a case study. Our formative study revealed ineffective and deceptive designs at all levels of a ToS. To make service contracts approachable, we designed and evaluated an intelligent reading interface: TermSight. Participants reported that TermSight reduced the difficulty of engaging with ToS and increased their willingness to do so. Additionally, features of TermSight enabled participants the ability to surface and approach relevant information at all levels of contract reading. Taken together, TermSight presents one avenue for making legal contracts more approachable to the general public. However, facilitating a deeper understanding of ToS remains an open challenge due to the overwhelming amount of information in ToS.
Appendix A Iterative Design Study
A total of 8 participants were recruited through Prolific to evaluate an early prototype of TermSight. This prototype closely resembled the current version of TermSight but lacked a document preview when hovering over the Power Meter. Additionally, the prototype used three saturation levels for relevance, resulting in a total of nine colors (3 hues for power x 3 saturation levels for relevance). Participants could also toggle between two layouts for the Summary Snippets: condensed layout and in-context layout. The condensed layout matched the current design shown in Figure  1. In contrast, the in-context layout placed each Summary Snippet directly next to its corresponding Information Snippet. After a brief interface tutorial, participants were given 10 minutes to read a ToS using the interface, followed by a semi-structured interview about their experience. Each session lasted for 40 minutes. Below, we present the main findings and changes made.